OpenStudy (anonymous):

Can someone explain why sinx/x^2 dx = -sinx/x

OpenStudy (turingtest):

? is there an integral sign missing?

OpenStudy (turingtest):

\[\frac{\sin x}{x^2}dx=\frac{-\sin x}x\]???

OpenStudy (anonymous):

yeah.

OpenStudy (anonymous):

still dont understand why thats the answer we get though.